OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomores tickets totaled $410. How many sophomores bought tickets? My answer is 120 Am I right?

OpenStudy (anonymous):

you're wrong 2x+3y=410 x+y=150 -> -2x-2y=-300 2x+3y-2x-2y=410-300 y=110 do you understand?

OpenStudy (anonymous):

Ahh yes! I see where I messed up, thank you! :)

OpenStudy (anonymous):

welcome